OpenStudy (anonymous):

A triangular prism has a volume of 10 cm3. If each dimension of the prism is multiplied by four, what is the new volume of the prism?

OpenStudy (anonymous):

320 cm3 80 cm3 160 cm3 640 cm3

OpenStudy (anonymous):

v= 0.5 * (b*h*l), each dimension is increased by 4 units v= 4* 4*4*0.5(bhl)= 32bhl

OpenStudy (anonymous):

i havent given the answer just expalined a general case of increasing each dimension by 4